if n is a positive integer,then (6*10^-n)+(1*10^-n) must equal?
Found 2 solutions by jim_thompson5910, RAY100:
Answer by jim_thompson5910(35256)
Answer by RAY100(1637)

6*10^-n + 1* 10^-n
.
6 / 10^n + 1/ 10^n = 7/10^n = 7*10^-n